### v1

# Tokenization is Sensitive to Language Variation

Abstract
Variation in language is ubiquitous and often systematically linked to regional, social, and contextual factors. Tokenizers split texts into smaller units and might behave differently for less common linguistic forms. This might affect downstream LLM performance differently on two types of tasks: Tasks where the model should be robust to language variation (e.g., for semantic tasks like NLI, labels do not depend on whether a text uses British or American spelling) and tasks where the model should be sensitive to language variation (e.g., for form-based tasks like authorship verification, labels depend on whether a text uses British or American spelling). We pre-train BERT base models for the popular Byte-Pair Encoding algorithm to investigate how key algorithmic design choices impact downstream models’ performances: fitting corpus, pre-tokenizer and vocabulary size. We find that the best tokenizer varies on the two task types—with the pre-tokenizer having the biggest impact on performance. Further, we introduce a new approach to estimate tokenizer impact on downstream LLM performance, showing significant improvement over techniques like Rényi efficiency. We encourage more work on language variation and its relation to tokenizers and thus LLM performance.
Tokenization is Sensitive to Language Variation
Anna Wegmann1, Dong Nguyen1 and David Jurgens2 1Utrecht University, Utrecht, The Netherlands 2University of Michigan, Ann Arbor, United States {a.m.wegmann, d.p.nguyen}@uu.nl, jurgens@umich.edu
[…]
Figure 1: Different types of tasks might profit from different tokenization settings. We investigate whether the same tokenizer performs equally well for tasks that require robustness to language variation (i.e., semantic-focused tasks like NLI) and for tasks that require sensitivity to language variation (i.e., form-focused tasks like authorship verification). Intuitively, one needs to look at semantic signals (e.g., “guitar” and “instrument”) for NLI and at form- or style-based signals (e.g., “learned” and “learnt”) for Authorship Verification.
Example Label
GLUE How do I hire an ethical hacker? Where can I find ethical hackers? duplicate
-typo Wher’re cacn I fien ethical hackprs? Hjw fo I hie an ehtical hacker? duplicate
-dialect How do I hire me an ethical hacker? Where can I find me ethical hackers? duplicate
(a) Tasks requiring robustness to language variation.
Dataset Task
AV Authorship Verification
PAN Author Change Classification
CORE Register Classification
NUCLE Error Classification
Dialect Dialect Classification
(b) Tasks requiring sensitivity to language variation.
Table 1: Evaluation Tasks. We evaluate tokenizers on tasks that require (a) robustness and (b) sensitivity to language variation. For (a), we use GLUE tasks, spelling perturbations using Wang et al. (2021) and dialect perturbations using Ziems et al. (2023). Examples are taken from QQP. For (b) we use a newly compiled selection of tasks containing our own authorship verification dataset, an author change classification dataset Bevendorff et al. (2024), a register classification dataset Laippala et al. (2023), a grammatical error correction dataset Dahlmeier et al. (2013) and a dialect classification dataset generated with Ziems et al. (2023).
1 Introduction
Variation in language is ubiquitous, manifesting in forms such as lexical variation (e.g., vocabulary choice like “lift” vs. “elevator”), spelling variation (e.g., spelling “u” instead of “you”), and syntactic variation (e.g., word order like “I gave her the ball.” and “I gave the ball to her.”). Such variation is often not random, but systematically linked to regional, social, and contextual factors Nguyen et al. (2016); Coupland (2007); Eckert (2012).
Tokenizers break up input strings and determine the tokens that are fed into language models. Tokenizers build their vocabulary based on a fitting corpus and have a given vocabulary size. A pre-tokenizer can prioritize or prevent the creation of certain tokens. Depending on these settings, tokenizers might behave differently for linguistic forms that are less common (see §3). For example, spelling variants of common words might be split into more tokens than their standard spelling Matthews et al. (2024), e.g., “doing” might be one token, while “doin” might be split into “do” and “in”. As a result, LLMs might need to learn to recompose such spellings to represent their meaning, suffering in performance compared to representing words as one token. In contrast, smaller tokens could be useful for recognizing form-based patterns (e.g., distinguish the “in” from the “ing” ending). We ask RQ1: Do the same tokenizer settings (i.e., fitting corpus, vocabulary size and pre-tokenizer) perform well for two types of downstream tasks: Tasks whose gold labels are robust to language variation (e.g., tasks that involve analyzing the semantic meaning of texts like NLI) and tasks whose gold labels are sensitive to language variation (e.g., tasks centered around the form, style or language variety of texts like authorship verification and dialect classification)? For example, for semantic tasks like NLI, models should perform equally well, regardless of whether the text is written in Standard American English or African American English. In contrast, for form-based tasks like authorship verification (i.e., are two texts written by the same author?), models should be able to distinguish British and American spelling. Intuitively a model needs to use different signals (see Figure 1) to solve tasks requiring robustness and tasks requiring sensitivity to language variation, and might thus profit from different tokenizer settings. We test this by pre-training BERT base models Devlin et al. (2019) with different tokenizer settings of the de-facto standard tokenizer: Byte-Pair Encoding or BPE Sennrich et al. (2016).
The evaluation of tokenizers with fast proxy measures, usually based on the token distribution in a reference corpus, offers a cheaper alternative to training LLMs for comparing different tokenizers. However, common proxy measures do not consistently achieve a high correlation with downstream performance Zouhar et al. (2023); Cognetta et al. (2024); Schmidt et al. (2024). One reason could be that such measures are task-agnostic. We therefore investigate RQ2: Do simple machine learning models like logistic regression with feature sets based on the different tokenizer vocabularies predict performance better than measures based only on the token distribution of the downstream corpus?
Contributions. We make the following three contributions: (1) We provide our code and a selection of English tasks (see Table 1) to evaluate models on tasks requiring sensitivity and robustness to language variation. [^1] GitHub link will be released with the peer-reviewed publication. (2) We find that tokenizer impact on downstream performance vary depending on whether a task requires robustness or sensitivity to language variation—with the pre-tokenizer having the biggest influence on downstream performance. (3) Further, we find that logistic regression performance has a higher correlation with BERT downstream performance than metrics like Corpus Token Count or Rényi efficiency Zouhar et al. (2023). We provide practical suggestions to evaluate and build better tokenizers. With this work, we hope to encourage more work on language variation and its relation to tokenizers to build better, fairer and more robust LLMs from the ground up.
2 Related Work
Next to Byte-Pair Encoding Sennrich et al. (2016), there exist several other subword (and other) tokenizers Mielke et al. (2021). For example, character and byte-based tokenizers have been argued to be more robust to spelling variations Mielke et al. (2021); Libovický et al. (2022); Xue et al. (2022). In this study we focus on BPE as it has been the most common tokenization algorithm in recent LLMs (e.g., Llama 3, Mixtral, DeepSeek v3 and GPT4).
There is no agreed-upon standard to evaluate tokenizers. Tokenizers have been evaluated intrinsically (i.e., without training LLMs) and extrinsically (i.e., considering the performance of LLMs pre-trained with the considered tokenizer). Common intrinsic methods include: the average number of subwords produced per word and correlated measures like Corpus Token Count Rust et al. (2021); Scao et al. (2022); Ali et al. (2024); Gallé (2019); Schmidt et al. (2024), information theoretic measures like Shannon Entropy or Rényi efficiency Zouhar et al. (2023) and morphological alignment Gow-Smith et al. (2022); Uzan et al. (2024). Extrinsic measures include LLM perplexity Shliazhko et al. (2024); Zevallos and Bel (2023); Gowda and May (2020), cross-entropy loss Rajaraman et al. (2024), computational training cost Ali et al. (2024) and downstream task performances Schmidt et al. (2024); Ali et al. (2024). It is computationally infeasible to train SOTA LLMs end-to-end for each version of the tokenizer one wants to evaluate. Recent work tackled this by training “smaller” generative language models with 350M–2.5B parameters for each tokenizer Schmidt et al. (2024); Ali et al. (2024). However, even such smaller models still can take several days to train. We measure downstream task performance on models with 110M parameters taking less than 15 GPU hours to train per model.
Tokenizer algorithms and settings can affect a LLM’s performance, for example, on tasks including numbers like arithmetic Thawani et al. (2021); Wallace et al. (2019), on tasks including domain specific vocabulary or jargon like coding or medicine Gu et al. (2021); Dehaerne et al. (2022); Zan et al. (2023), on different scripts and languages Petrov et al. (2023); Rust et al. (2021); Limisiewicz et al. (2023); Ahia et al. (2023); Velayuthan and Sarveswaran (2025) and when translating between languages Gallé (2019); Libovický et al. (2022); Zhang et al. (2022). To the best of our knowledge monolingual tokenizers have been underinvestigated in relation to language variation. Monolingual tokenizers and tokenizer settings have recently been investigated on broader selections of NLU tasks Schmidt et al. (2024); Ali et al. (2024), presumably with the underlying assumption that for a given language like English, there exists a best tokenizer for most, if not all, tasks. We investigate this assumption for two systematically different types of tasks: tasks requiring robustness and tasks requiring sensitivity to language variation.
3 Tokenizer Settings
Tokenizers break up texts into smaller units. These units are fed into the language model as input. We investigate different variations of the most popular tokenization algorithm: Byte-Pair Encoding or BPE Sennrich et al. (2016). BPE is a subword tokenization algorithm (i.e., it breaks rare words down into subword units, bottoming out in bytes in the worst case). The vocabulary is built iteratively: It starts out with a base vocabulary of distinct bytes and merges them based on token frequency in the fitting corpus until the desired vocabulary size is reached. We vary the algorithm on three parameters: (1) the vocabulary size, (2) the pre-tokenizer and (3) the fitting corpus.
3.1 Vocabulary Size
Common vocabulary sizes range from 30k to 64k in monolingual models, to about 128k (Llama 3) and 200k (GPT 4o) in recent multilingual models. To the best of our knowledge, previous work on tokenizer vocabulary size mostly tested mid-range vocabulary sizes between 32k and 100k on NLU tasks, finding only small differences in performance between vocabulary sizes Ali et al. (2024); Schmidt et al. (2024). However, vocabulary size might still play a role when dealing with language variation. When using mid-range vocabulary sizes, less common words or spelling variants might not be represented with one token anymore, influencing a word’s position in the embedding space Matthews et al. (2024). Character-based models like ByT5 show better robustness to spelling variation than subword models with larger vocabulary sizes Libovický et al. (2022); Xue et al. (2022); Tay et al. (2021). Similarly, tokenizers using low-range (in the extreme case character-level) vocabulary sizes might be more robust to spelling variation, as, for example, swapping of characters will not drastically alter the character-based segmentation of a word. We experiment with the following vocabulary sizes: 500 and 4k (low-range vocabulary size), 32k and 64k (mid-range vocabulary size) and 128k (high-range vocabulary size).
3.2 Pre-Tokenizer
Pre-tokenizers split the fitting corpus into word-like units Mielke et al. (2021), so-called “pre-tokens” (e.g., “I’m fine.” $\rightarrow$ “I”, “’m”, “fine”, “.”) which are then further tokenized independently. That is, in building its vocabulary, the BPE algorithm cannot merge tokens that cross pre-token boundaries (e.g., “I’m”). Pre-tokenizers are usually expressed with regular expressions that include hard-coded knowledge about a language or script Velayuthan and Sarveswaran (2025), e.g., whitespaces separate words in Latin script and “’m” is a contraction in English. The most common English pre-tokenizers are more or less elaborate variations of splitting texts based on whitespaces. For example, GPT-2 increasingly separates different Unicode Character Categories (e.g., letters, numbers and punctuation) but leaves single leading whitespaces attached to words Radford et al. (2019).
Pre-tokenizers influence compression efficiency Radford et al. (2019), NLU downstream Schmidt et al. (2024) and arithmetic performance.222right-to-left pre-tokenization of numbers improved arithmetic for Llama3, see bluesky. However, pre-tokenizers have not been investigated for tasks that require sensitivity to language variation. Pre-tokenizers might play a role for words that merge Unicode Character Categories, for example, when letters are substituted with numbers (e.g., “2day” or “c000l”, see Eger et al. (2019) for more). If a pre-tokenizer generally splits off numbers and letters, such words can never become part of the vocabulary. Further, not splitting on whitespace might better represent syntactic variation by including frequent phrases like “of the”. Frequent compositional phrases are processed faster by humans Arnon and Snider (2010), and building tokenizers that align closer to such cognitive processes (e.g., Yang et al., 2020) might also be beneficial for semantic tasks.
We compare five different pre-tokenizers: (a) not using a pre-tokenizer ( N O ), (b) isolating whitespaces ( W S ), (c) leaving single whitespaces attached to words (_ W S ), (d) the L L A M A 3 pre-tokenizer which splits off Unicode Character Categories (e.g., punctuation) but leaves one leading non-letter character attached to letters (e.g., “’m”) and (e) the widely used G P T 2 pre-tokenizer which separates Unicode Character Categories the most. See Appendix A.2 for details.
3.3 Fitting Corpus
Sub-word tokenizers construct their vocabulary based on a fitting corpus, adding tokens based on the distribution of tokens in that corpus. Lexical, morphological and spelling variation are all intuitively connected to the fitting corpus. If a fitting corpus does not show a variation, it can never be part of a single token in the vocabulary.
We sampled fitting corpora with a size of approximately 1.5 billion words. PubMed was randomly sampled from The Pile’s PubMed Abstracts Gao et al. (2020). The PubMed Abstracts consist of 30M abstracts from biomedical articles. Wikipedia was randomly sampled from Wikipedia articles from a snapshot from June 1st, 2023 after the plain text of articles was extracted.333using https://github.com/LorcanJConnolly/WikiTextExtractor. Twitter was sampled from the Decahose Twitter stream [^4] Decahose provided access to 10% of real time tweets sampled by Twitter. throughout the year 2021, with queries on almost every day of the year 2021. We only select English tweets based on Twitter’s internal language identification system. We exclude retweets. Miscellaneous was sampled from a variety of domains with no overlap with the other fitting corpora. It includes Reddit, literature sources (fanfictions and books), news articles and comments, question answering websites, reviews, mails, transcripts, blogs, Common Crawl, scientific articles, code and mathematical formulas. See details in Appendix A.1.
We expect the fitting corpora to differ in the lexical, syntactic and lexical variation they exhibit. For example, PubMed probably contains less spelling and lexical variation than Twitter. Intuitively, a tokenizer constructed on PubMed should thus be less capable in representing stylistic variation than one constructed on Twitter. However, it remains unclear how important the fitting corpus composition is. Zhang et al. (2022) investigate different compositions of the fitting corpus in the multilingual setting. They find a surprising robustness to language imbalance in the fitting corpus for languages sharing the same script.
4 Evaluation Tasks
We compare tokenizers on classification tasks that require robustness to language variation (§4.1) and tasks that require sensitivity to language variation (§4.2). See an overview in Table 1. These two types of tasks need to make use of more semantic and form-based signals respectively and might have different requirements for a tokenizer (cf. Figure 1).
4.1 Tasks Robust to Language Variation
First, we evaluate tokenizers on semantic tasks where the gold label is robust to language variation. We use GLUE (Wang et al., 2018) as is a common NLP benchmark and was used to evaluate BERT when it was originally introduced Devlin et al. (2019). We compare tokenizers on the following four GLUE tasks SST-2 (sentiment classification), QQP (paraphrase classification), MNLI and QNLI (NLI tasks). For details on task selection see Appendix LABEL:App:NLU. Ideally, a robust tokenizer performs consistently across all versions of GLUE: the original, primarily written in Standard American English, the spelling-transformed GLUE-typo and the dialect-transformed GLUE-dialect.
GLUE-typo. We use textflint Wang et al. (2021) to introduce simulated typos and spelling errors to our tasks, similar to Libovický et al. (2022). It uses random character swapping and a list of common spelling errors.
GLUE-dialect. We use Multi-VALUE Ziems et al. (2023) to introduce simulated dialectal variation to our GLUE tasks. Multi-VALUE makes use of 189 dialectal perturbation rules (Ziems et al., 2023) operationalized based off of eWAVE Kortmann et al. (2020). For each example in the GLUE tasks, we randomly choose between transformations to Appalachian English, Chicano English, Colloquial Singapore English, Indian English and Urban African American English dialects.
4.2 Tasks Sensitive to Language Variation
Second, we evaluate tokenizers on their downstream performance on tasks that are sensitive to language variation. The gold label for the task should use stylistic or form-based signals.
AV. Past work has found authorship verification (i.e., whether two texts were written by the same author) highly sensitive to tokenization, with significant gaps between the BERT and RoBERTa tokenizers (Zhu and Jurgens, 2021). Therefore, we curate a new authorship verification dataset of 40.8k train, 2.5k dev and 4.8k test pairs of texts from different domains, similar in distribution to the Miscellaneous corpus (c.f. §3). Labels are balanced in the test set. See details in Appendix LABEL:app:sadiri.
PAN. We use the PAN 2024 Multi-Author Writing Style Analysis task to predict whether an author shift occurs between two consecutive paragraphs extracted from Reddit Bevendorff et al. (2024). Specifically, we use the ‘hard’ task, where paragraphs are about the same topic. We sample such that labels are balanced. This results in a training set of 18k and a dev set of 4k instances.
NUCLE. We use the NUCLE 3.3 corpus Dahlmeier et al. (2013) for multi-label classification of the errors that were made by English learners in a given sentence. NUCLE was annotated by professional English instructors for 27 error types (e.g., verb tense or article use). It includes 22k unique sentences with errors. We add a sample of 5k error-free sentences from the same dataset. We split it into a train (80%) and dev (10%) set.
CORE. We use the Corpus of Online Registers of English Laippala et al. (2023) for register classification. Register is one of the most important factors associated with linguistic variation Biber (2012). We use 8 main register labels (e.g., spoken or informational description) for multi-class prediction. To increase the occurence of rarer labels, we split long texts and reach a train size of 30k and a dev size of 5k. See Appendix LABEL:app:sadiri for details.
Dialect. We randomly sample 60k instances from GLUE-dialect (§4.1) and the original GLUE task, to create a dialect classification task for five dialects and Standard American English in the original GLUE texts. We use 50k texts for the train and 5k for the dev set.
5 Modeling
For each investigated tokenizer (cf. §3), we pre-train one BERT model. We use encoder instead of decoder models, as encoder models tend to reach higher performance for classification tasks for low parameter settings. This allows us to train models using fewer GPU hours.
Pretraining BERT Models. We experimented with pre-training BERT models using a token count T that was exponentially bigger than the number of parameters P, similar to the original BERT papers Devlin et al. (2019); Turc et al. (2019). However, we find that for the same compute, using a bigger model size P and less tokens T improves the training loss. Chinchilla’s scaling law might also hold for smaller encoder models, specifically optimal parameter count could scale with the token size for a fixed compute $P_{OPT}\approx T^{23/27}$ Hoffmann et al. (2024). For the remainder of this work, we use the base BERT model architecture with 110M parameters, initialize all weights randomly and pre-train on 750M tokens sampled in sequences of 512 from the Miscellaneous corpus (§3.3) and use a batch size of 32 and 45k steps. For further details and hyperparameters, see Appendix LABEL:app:pretraining.
Fine-Tuning BERT. Unless otherwise specified, we used 3 epochs, a max sequence length of 128, a batch size of 32 and a learning rate of 2e-5 to perform the classification tasks. We evaluate on the dev set for GLUE. For comparability, we use the same setup for tasks requiring sensitivity to language variation. Only for the authorship verification task we use a contrastive training setup, then use the dev set to find an optimal cosine similarity threshold and calculate accuracy on the test set.
Figure 2: Pairwise significance testing of models. We use pairwise McNemar (1947)’s test to test how differently models trained with different tokenizers classify the tasks requiring robustness to language variation (first row) and tasks requiring sensitivity to language variation (second row). Tokenizers are sorted by mean performance. Blue colors show statistical significance, while red colors are above the 0.05 threshold.
6 Results
We show the performance of fine-tuned BERT models on tasks that require robustness to language variation in Table 2 and tasks that require sensitivity to language variation in Table 3. Note that the performance differences tend to be relatively small, which is consistent with previous work comparing different tokenizer algorithms on downstream tasks Ali et al. (2024); Schmidt et al. (2024). We compute the pairwise McNemar, 1947’s test for the pre-tokenizer, fitting corpus and vocabulary size settings, see Figure 2. We use the Bonferroni correction Dunn (1961) for our total of 26 tests.
RQ1: Tokenizer settings perform differently on tasks robust and sensitive to language variation. Tasks sensitive to language variation profit from tokenizers that encode more variation through: (i) a larger vocabulary size (§6.3) and (ii) pre-tokenizers that combine different Unicode Character Categories like punctuation and letters (§6.1).
Model orig +typo +dialect AVG
Corpus PMed 80.8 68.9 78.6 76.1
Wiki 80.7 68.3 79.3 76.1
Twitter 81.1 68.7 78.7 76.2
Pre-Tok
N
O
73.4 61.7 69.8 68.3
W
S
80.3 67.7 79.0 75.7
_
W
S
80.9 68.9 78.8 76.2
L
L
A
M
A
3
81.1 68.0 79.0 76.0
G
P
T
2
81.7 68.4 79.7 76.6
Size 500 78.5 71.6 76.6 75.6
4k 80.1 69.8 78.2 76.0
32k 81.7 68.4 79.7 76.6
64k 81.2 68.3 79.4 76.3
128k 81.4 67.4 79.6 76.1
Table 2: Performance on Tasks Robust to Language Variation. We display BERT performances, averaged on the original four GLUE tasks and their perturbations using spelling mistakes (+typo) and dialectal transformations (+dialect). We boldface the best performances. For the averaging column (AVG), italics indicate tokenizers with no statistically significant difference from the best-performing tokenizer.
Model AV (acc) $\uparrow$ PAN (acc) $\uparrow$ CORE (acc) $\uparrow$ NUCLE (F1) $\uparrow$ DIALECT (F1) $\uparrow$ Mean
Corpus PubMed 81.7 65.6 56.5 23.0 88.0 63.0
Wikipedia 81.8 64.8 54.9 23.7 88.8 62.8
Twitter 83.4 66.6 56.7 22.1 88.1 63.4
Pre-Tok.
N
O
82.2 58.5 51.8 16.0 77.5 57.2
W
S
74.2 65.9 55.0 22.9 88.9 61.4
_
W
S
82.1 65.8 57.5 22.3 88.4 63.2
L
L
A
M
A
3
82.7 67.0 57.3 21.6 88.6 63.4
G
P
T
2
82.7 64.9 57.9 21.7 88.8 63.2
Size 500 79.2 62.7 51.6 14.1 58.7 58.7
4k 81.9 67.8 56.2 15.2 88.2 61.9
32k 82.7 64.9 57.9 21.7 88.8 63.2
64k 82.8 66.3 56.4 23.6 88.8 63.6
128k 82.7 67.1 54.3 22.8 88.7 63.1
Table 3: Performance on Tasks Sensitive to Language Variation. We display the BERT performance on our Authorship Verification (AV), the PAN, the CORE, and the multi-Dialect dataset. According to mean performance values, the best fitting corpus is Twitter, the best pre-tokenizer is
GPT
2 and the best model size is 64k. For the mean column, italics indicate tokenizers with no statistically significant difference from the best-performing tokenizer.
6.1 Pre-Tokenizers
Pre-tokenizers have the greatest influence on performance. For both task types, the range of performance values is largest for pre-tokenizers, second largest for vocabulary size and smallest for fitting corpus. This is surprising as, to the best of our knowledge, pre-tokenizers have received the least attention in previous work.
Pre-tokenizers improve performance. Using pre-tokenizers improves performance over using no pre-tokenizer. Combining leading whitespaces with letters (_ W S ) generally improved performance over isolating whitespaces ( W S ). A reason could be that splitting off whitespaces increases, in the worst case doubles, the number of input tokens to the model forcing it to reason over longer sequences Radford et al. (2019).
Combining Unicode Character Categories is more useful for tasks requiring sensitivity to language variation. For both task types, L L A M A 3 and G P T 2 are among the best performing pre-tokenizers. One of their main differences to other pre-tokenizers is that the combination of different Unicode Character Categories (e.g., punctuation and letters) within the same token is less often allowed. Overall, G P T 2 allows for the fewest combinations of Unicode Character Categories. By preventing combinations of categories like punctuation and letters, the vocabulary of G P T 2 can include a broader range of tokens that only consist of letters, e.g., “_queer” in Table 4. This could explain the better performance of G P T 2 on tasks that require robustness to language variation. In contrast, L L A M A 3 performed better on tasks requiring sensitivity to language variation. Allowing the mixing of one initial punctuation mark with letters, e.g., including “’all” in Table 4, seems to help L L A M A 3 solve PAN.
6.2 Fitting Corpus
Twitter performs best on tasks requiring sensitivity to language variation. This aligns with our expectation that the Twitter corpus include more spelling variation and a larger set of language varieties than PubMed and Wikipedia.
Twitter performs surprisingly well on tasks requiring robustness to language variation. Interestingly, Twitter performs indistinguishably from other fitting corpora on tasks that require robustness to language variation. Originally, we expected a more standardized corpus like Wikipedia to perform better, as it should lead to less spelling variation and thus more “full words” in the vocabulary (e.g., “precursor” in Table 4).
6.3 Vocabulary Size
A larger vocabulary size might be useful for tasks requiring sensitivity to language variation. Performance peaked at 32k for tasks requiring robustness to language variation and at 64k for tasks requiring sensitivity to language variation. It seems that a vocabulary size of 32k is sufficient to learn common standard forms and more tokens are needed to include style and form variations. For example, “jumper” is only included with vocabulary size 64k, see Table 4. Note that optimal vocabulary sizes might scale with model size Tao et al. (2024).
Smaller vocabulary is more robust to spelling variation. The vocabulary size of 500 consistently performs worse than larger vocabulary sizes. It seems that BERT struggles more with long input sequences and learning to compose words for small vocabulary sizes. An exception is the GLUE-Typo task, where the tokenizer of size 500 performs best. This is consistent with character-based tokenizers being more robust to spelling variation (cf. §3.1).
7 Pre-evaluating Downstream Tokenizer Impact
An popular alternative to training several LLMs to evaluate tokenizers is using fast intrinsic measures on the token distribution on the downstream task corpus. However, common intrinsic measures are not consistently correlated with downstream performance Rust et al. (2021); Zouhar et al. (2023); Schmidt et al. (2024). Further, they do not make use of task labels and can thus be considered task agnostic. Imagine a corpus on which you want to perform both, a task requiring robustness to language variation and a task requiring sensitivity to language variation (cf. Figure 1). Task-agnostic intrinsic measures like Rényi efficiency and Corpus Token Count will always predict the same performance for both tasks. However, we found that the type of task can have an influence on what tokenizer settings work better (cf. §6).
We experiment with a task-dependent measure: logistic regression with task labels as the dependent variable. Specifically, we use a tokenizer’s vocabulary as the feature set with a bag-of-token approach. For tasks with two texts as inputs, token combinations between sentence pairs serve as features. For multi-class and multi-label tasks, we use one-vs-rest models. See Appendix LABEL:app:log-regression for details.
Group Choice Examples of tokens
Corpus Wiki _queer _precursor _l ma o
Twitter _queer _prec urs or _lmao
PubMed _qu e er _precursor _l ma o
Pre-Tok
N
O
_ y’all_ que er_ b 4
W
S
_ y’all _ que er _ b4
_
W
S
_y’all _que er _b 4
L
L
A
M
A
3
_y ’all _que er _b 4
G
P
T
2
_y ’ all _queer _b 4
Size 500 _t e h _l g b t q i a + _j um p er
4k _te h _l g bt q ia + _j um per
32k _te h _l g bt q ia + _j umper
64k _teh _lg bt q ia + _jumper
128k _teh _lgbt q ia + _jumper
Table 4: Examples of Vocabulary Differences. We display examples of sequences tokenized by tokenizers with varying fitting corpora, pre-tokenizers and vocabulary size. We represent whitespaces within tokens as _. The domain of the fitting corpus affects the set of unique words in the vocabulary (e.g., “queer” is part of Twitter but not PubMed). The size determines the number of “rarer” words in the fitting corpus that are added to the vocabulary (e.g., the British variant “jumper” of “sweater” is only added with size 64k). The pre-tokenizer determines what types of words can be part of the vocabulary (e.g., “b4” can not be part of
G
P
T
2 and
L
L
A
M
A
3). Note that allowing less Unicode Character Category combinations (e.g., numbers and letters) increases the broadness of the vocabulary (e.g.,
G
P
T
2 allows for less combinations than other pre-tokenizers and is the only one including “queer” in its vocabulary).
7.1 Results
We display the Pearson correlation of Rényi efficiency, Corpus Token Count and logistic regression performances with the fine-tuned BERT performances in Table 5. Similar to Schmidt et al. (2024) we only find a light negative correlation with Rényi efficiency.
Corpus token count is conversely correlated with downstream performance, showing a negative correlation for tasks requiring robustness and a positive correlation for tasks requiring sensitivity to language variation. Corpus Token Count and correlated measures are often used to assess the compression quality of a tokenizer on a reference corpus or downstream task. A higher value entails more tokens in the corpus and thus worse compression. Better compression is commonly believed to be a sign of a better tokenizer, and has often been thought to be correlated with better downstream performance Rust et al. (2021); Ali et al. (2024); Velayuthan and Sarveswaran (2025). We show that the correlation is flipped for tasks that are robust and sensitive to language variation highlighting the difference in tokenizer requirements for the two task types.
RQ2: Logistic regression correlates with downstream performance. Among the three considered measures, we find the highest correlation between logistic regression performances and BERT performances. Additionally, logistic regression is similarly correlated for both task types and can compare tokenizers of varying vocabulary sizes, which is not possible with Rényi efficiency and Corpus Token Count (for details refer to Appendix LABEL:app:intrinsic_vocab-size). Note that we by no means question the usefulness of measures like Corpus Token Count or Rényi efficiency for assessing different tokenizers and what they do. However, they might be less suited to estimate downstream performance without additional modifications.
Measure Robust Sensitive
Rényi Efficency -.20 -.03
Corpus Token Count -.45 .37
logistic regression .86 .86
Table 5: Correlation of Intrinsic Measures with BERT Performances. Correlations between the Rényi Efficency ($\alpha=2.5$), Corpus Token Count, and logistic regression predictions with the performance of the finetuned BERT models on the tasks robust to language variation and the tasks sensitive to language variation respectively. Logistic regression has the highest correlation values across all tasks. A higher Corpus Token Count is negatively correlated with performance on tasks robust to language variation and positively correlated with tasks sensitive to language variation.
8 Conclusion
In this work, we investigated tokenizer settings for tasks that require robustness and tasks that require sensitivity to language variation. BPE settings perform differently on the two task types. We make three practical suggestions for selecting tokenizer settings: (1) Pay the most attention to the pre-tokenizer. It influences how Unicode Character Categories can be combined (e.g., “b4” in Table 4), and how many different words are ultimately part of the vocabulary (e.g., “_queer” in Table 4). (2) Choose a bigger vocabulary size for settings that require sensitivity to language variation. (3) Use a small machine learning classifier to test the effect of different tokenizers on tasks in your domain of interest. In the future, it might be helpful to make intrinsic measures sensitive to tasks when used to estimate model performance. With this work, we hope to encourage further work on language variation and its relation to tokenizers to ultimately build better, fairer and more robust LLMs.
Limitations
Note that our taxonomy of tasks relies on splitting tasks into semantic-focused tasks (i.e., considering what is said) and form-based tasks (i.e., considering how it is said). However, a strict distinction is challenging since most tasks are best solved using a mixture of typical content and typical form signals. For example, recognizing the lengthening of words (e.g., “cooool”) can be helpful for the semantic task of sentiment classification Brody and Diakopoulos (2011) and content information (e.g., the topic) can improve authorship verification performance Wegmann et al. (2022). Still, this distinction clarifies the main signals that our different task types should rely on and enables us to compare tokenizer settings on these two different types of tasks.
We evaluate tokenizers on 110M parameter encoder models as encoder models tend to reach higher performance for classification tasks in low parameter settings. However, we risk not accounting for emergent capabilities of popular larger generative language models. For example, the number of context tokens for recent models are generally much larger than the original 512 tokens of BERT base.
We find that logistic regression correlates with downstream performance on classification tasks. For more complex tasks like regression or generation tasks, logistic regression might not be applicable. However, we still see it as a great benefit that logistic regression can enable us to quickly test tokenizers on a variety of different tasks and make a more informed decision on what tokenizers to test more rigorously.
For tasks requiring robustness to language variation, many more tasks could have been included. Future work could investigate tasks such as question answering or arithmetic tasks. For tasks requiring sensitivity to language variation, we originally planned to include more classification tasks that are tied to well-established factors of language variation, e.g., age prediction and geographic location prediction Nguyen et al. (2016). However, we repeatedly encountered difficulties with cleanly separating semantic from form-based cues. For age prediction and location prediction specifically, logistic regression models made extensive use of content words. Based on the coefficients of our logistic regression models, we excluded tasks that we expected to mainly require sensitivity to language variation, but where the models primarily relied on content words.
The tokenizer settings seem to have different effects when the pre-training corpus and fitting corpus differ. The pre-training corpus might influence in how far a model can leverage the diversity encoded in the tokenizer. See Appendix LABEL:app:webbook-pretraining for further experiments. Note, however, that recent work suggests to only use vocabulary that appears in the training corpus Land and Bartolo (2024) – which is unlikely when different corpora are used for model training and tokenizer fitting.
Ethical Considerations
Our pre-training and fitting corpora are largely based on publicly shared and accessible datasets from popular online forums and web pages (e.g., Wikipedia, Reddit, Common Crawl, …). Unfortunately, these datasets were mostly collected without explicit consent from users and might lead to (among others) privacy concerns. Individuals might be identifiable from their written texts. However, we hope that the risks of reusing already published datasets are minimal. We collected tweets from Twitter in 2021 using Twitter’s official API, but opt to not share them publicly. While we aim to include different language varieties in our datasets, they might not be representative of English language use across different social groups. For example, we expect a skew towards straight, white, American, young and male authors. We caution against using our datasets and tasks for general claims about broad selection of language varieties. We confirm to have read and that we abide by the ACL Code of Ethics. Beside the mentioned ethical considerations, we do not foresee immediate risks of our work.
Acknowledgements
We thank the colleagues in the various NLP Groups at Utrecht University and the Blablablab at the University of Michigan and, specifically, Kees van Deemter, Antal van den Bosch, Yupei Du, Qixiang Fang, Melody Sepahpour-Fard, Shane Kaszefski Yaschuk, Elize Herrewijnen, Sanne Hoeken, Hugh Mee Wong, Jian Zhu and Pablo Mosteiro for, among others, feedback on writing and presentation. We thank colleagues at EMNLP 2024 for their insights into tokenization, and, specifically, Vilém Zouhar. We thank Laura Wegmann for feedback on writing.
Appendix A Tokenizer
In this section, we provide additional information on the tokenizer settings we investigate. See Table 6 for examples of unique tokens for each setting.
A.1 Fitting Corpora
See an overview of the fitting dataset sizes in Table 7. Usually, the fitting corpus for the tokenizer and the training corpus for the language model with that tokenizer are the same. As a result the size of the fitting corpus often varies as widely as the size of training datasets. We aim for 1.5 billion tokens for all fitting corpora. The MiniPile Kaddour (2023) dataset used for fitting in Schmidt et al. (2024) is of similar magnitude. We further display text examples for each dataset in Table 8. Dataset sizes vary in less than 1% of word count. The variance in word count is an artifact of dataset creation from several documents with lenient word count limits.
Miscellaneous. Miscellaneous consists of Reddit Baumgartner et al. (2020), literature sources (fanfictions from ao3555fanfictions until 2019 from Archive of Our Own https://archiveofourown.org/, downloaded from https://archive.org/download/AO3_story_dump_continuing in 2023, filtered for English language using AO3 tags. Dataset was removed but should be re-creatble using tools like https://github.com/nianeyna/ao3downloader. , Gao et al., 2020’s books before 1919 from the Gutenberg project), news articles and comments (Zellers et al., 2019’s realnews, 2020 NYTimes articles and comments [^6] https://www.kaggle.com/datasets/benjaminawd/new-york-times-articles-comments-2020 , minimum length filter of 250 was applied , Kolhatkar et al., 2020’s sfu-socc), question answering (Gao et al., 2020’s StackExchange), reviews (Hou et al., 2024’s Amazon and Wan and McAuley, 2018’s GoodReadsreviews), mails (Bevendorff et al., 2020’s Gmane), transcripts (YouTubeCommons [^7] https://huggingface.co/datasets/PleIAs/YouTube-Commons and Gao et al., 2020’s OpenSubtitles), blogs (Schler et al., 2006’s blogcorpus), raw text from webpages (Gao et al., 2020’s Common Crawl), science articles (Lo et al., 2020’s s2orc), code and mathematics (Gao et al., 2020’s GitHub and DeepMind Mathematics). See the share of different domains in Table 9.
A.2 Pre-Tokenizer
See the regular expressions defining the different considered pre-tokenizers in Table LABEL:app:pre-tokenizer. Differences affect mostly whitespace, contraction, punctuation and number handling.
name RegEx Example Text
[…]
ws